# Best Practices for Multi-Fidelity Bayesian Optimization in Materials and Molecular Research

### v2

@@ Abstract
Among Machine Learning (ML) techniques, Bayesian optimization (BO) has emerged as the go-to choice for optimizing the design of experiments in the chemical domain 1. Bayesian optimization, grounded in a probabilistic framework 2, consists of two main components: a probabilistic model that serves as a proxy for the experimental process being optimized, and a policy that governs the acquisition of new experimental data. For instance, a researcher seeking to maximize the yield of a given reaction would query the model to identify which experimental conditions to be tested to achieve the desired outcome. This methodology has shown success in diverse optimization tasks such as chemical reactions 1, 3, 4 and functional molecules 5. Iterating through this learning cycle has recently enabled the rapid identification of optimal conditions within extensive search spaces 6, 7.
@@
While canonical BO has recently been popularized among experimentalists, the experiment design may benefit if the practitioner can collect data at different degrees of reliability while paying a lower price. Additional experimental evidence that may be readily available can be integrated within the model representing the process to be optimized (e.g., low-precision experiments conducted with bench-top nuclear magnetic resonance can be integrated into more expensive, high-precision experiments 8). In classical experimental design, the inclusion of such information sources representing different reliabilities is referred to as multi-fidelity Bayesian Optimization (MFBO). Within this setting, a specific cost is assigned to each information source — hereafter defined as fidelity. The multi-fidelity probabilistic model learns the process of interest by extracting knowledge from data available at different fidelities and understanding their interplay. The policy for querying new experimental data at a given fidelity also takes into account the overall cost. By combining low-fidelity (LF) and high-fidelity (HF) points, the overall optimization cost can be reduced compared to the single-fidelity BO (SFBO). Figure 1 exemplifies how the iterative cycle of MFBO may reduce the overall cost, compared to the standard SFBO.
@@
Although MFBO has garnered interest in the past decade within the ML community, leading to a myriad of model definitions and acquisition policies 9, 10, 11, 12, researchers have only lately started to integrate them within their design methodologies. In the chemical domain, there has been recent interest in incorporating cost awareness into the BO loop13, 14. Regarding the specific multi-fidelity approach, several studies have successfully applied MFBO to the materials discovery domain 15, 16, 17, 18, 19, 20. However, there is a fundamental lack in the assessment of MFBO performance, and due to this, each work represents MFBO performance in different ways 21, 16, 19, 20. Although some metrics have been proposed 15, 19, the lack of clarity and unified criteria on how to assess the benefits of MFBO in the chemical domain has hindered its widespread adoption among the experimental community22. Several factors come into play when assessing its impact, yet the application of MFBO can also be detrimental to the overall optimization process 23. It has been also shown how in the long run the advantage of MFBO over SFBO can be lost 22. Therefore, it is crucial to unify and provide a reliable method to decide when MFBO is better than SFBO.
@@
In this work, we propose a series of guidelines on when to use MFBO within the experimental design pipeline. We conduct an exhaustive experimental investigation on both standard MFBO problems — ubiquitous in BO literature — as well as chemistry-based ones. Initially, we optimize two synthetic problems to assess the behavior of a multi-fidelity model in simulated black-box scenarios, detecting unfavorable situations where the application of MFBO does not offer an advantage over SFBO. We exhaustively scan MFBO experimental parameters (namely, cost ratio and informativeness of the LF source) to identify trends that indicate promising experimental settings for when to apply MFBO. Building on these guidelines, we then progress to chemistry-based problems, tackling three real challenges in molecular and materials optimization, where MFBO successfully outperforms SFBO. This work offers a comprehensive view and a reference guide for applying MFBO in the molecules and materials discovery domain.
@@ Methods
where f is the target black-box function and $\mathcal{X}$ is the space of all possible input candidates. Considering $\mathcal{D}=[(x_{1},y_{1}),(x_{2},y_{2}),...(x_{n},y_{n})]$ a dataset containing pairs of inputs-outputs $({x}_{n},y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ of our problem, we can use a surrogate model to learn the relationships between the data in the form $y_{i}=f({x}_{i})+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. The surrogate is commonly a Gaussian Process (GP), a non-parametric model that provides uncertainty quantification 24. GPs are usually defined as $f(x)\sim\mathcal{GP}(0,k({x},{x^{\prime}}))$, placing a zero-mean prior and a covariance given by a kernel function $cov[f({x}),f({x^{\prime}})]=k({x},{x^{\prime}})$ that measures the similarity between inputs. Given a dataset $\mathcal{D}$, the predictive mean and variance are given by the following expressions
@@
where ${k_{x}}=[k({x},{x_{1}}),...,k({x},{x_{n}})]^{\top}\in\mathbb{R}^{n}$, $K\in\mathbb{R}^{n\times n}=(k({x_{i}},{x_{j}}))_{1\leq l_{0},l_{1}\leq n}$ and ${y}=[y_{1},...,y_{n}]$. In order to optimize the target problem, after training the surrogate with the available data, a decision is made to select the next points by maximizing the acquisition function. Acquisition functions are heuristics that decide the next point to query by balancing exploration and exploitation. The next inputs are given by
$\operatorname*{argmax}_{{x}\in\mathcal{X}}\texttt{acqf}({x})$
where acqf is the acquisition function.
Our multi-fidelity BO extends the previous setup by considering an extra input parameter for the surrogate model, the fidelity l. This parameter indicates which source of information the other inputs correspond to (e.g., high fidelity, HF, or low fidelity, LF). In our case, we use a discrete fidelity setting, where $l\in[l_{0},l_{1},...l_{m}]$ and $l_{m}$ corresponds to the highest fidelity among the $m$ levels. In all cases, we limit the analysis to two levels of fidelity ($l_{0}$ and $l_{1}$ and $m=1$) , but the methodology can be extended to more levels. Additionally, these models assume that the cost only depends on the fidelity level $l$ and not on the search space $x$. We employed a modified GP proposed in 11 that extends the previously described surrogate modelling to the multi-fidelity space of $\mathbb{R}^{n+1}$ dimensions, chosen due to its popularity amongst MFBO applications11, 16, 23. The modified kernel is defined as $k(({x},l),({x^{\prime}},l^{\prime}))$, where fidelity is introduced by defining a separate kernel to model the input space, and a kernel to model the correlation and interaction between the different fidelity levels. The mathematical expression is:
@@
Given a dataset of inputs with their associated fidelities and outputs, the model can be trained using standard maximum log-likelihood optimization to obtain the optimal hyperparameters $\lambda_{i}$, $c$ and $\delta$. In the experiments, we set the $l$ values of the HF and LF levels to 1 and 0 respectively following previous studies that used this kernel in a binary setting16, 23.
@@
$\operatorname*{argmax}_{(x,l)\in\mathcal{X}\times[m]}\texttt{acqf}(x,l)$
In the experiments, we use two families of acquisition functions, with their single and multi-fidelity versions: Maximum Entropy Search (MES)12 and Expected Improvement (EI)25. We also initially tested Knowledge Gradient (KG)11, but we discarded it due to its computational expense.
In all cases, the experiments are run for 20 independent seeds, and the number of initial samples corresponds to 10% of the total optimization budget, sampled using a Latin hypercube strategy. In the MFBO case, the budget is split to sample 50% of HF and 50% of LF initial points (this proportion was selected based on preliminary experiments that showed how this ratio did not affect the final MFBO performance). In the synthetic functions, the total budget corresponds to 50 HF queries. In the chemistry and materials benchmarks, the total budget corresponds to 30 HF queries. Data collection is sequential. All the experiments are run using BOTorch 26 as a standard Bayesian Optimization framework.

### v3

@@ Introduction
Among Machine Learning (ML) techniques, Bayesian optimization (BO) has emerged as the go-to choice for optimizing the design of experiments in the chemical domain 1, 2, 3, 4. Bayesian optimization, grounded in a probabilistic framework 5, consists of two main components: a probabilistic model that serves as a proxy for the experimental process being optimized, and a policy that governs the acquisition of new experimental data. For instance, a researcher seeking to maximize the yield of a given reaction would query the model to identify which experimental conditions to be tested to achieve the desired outcome. This methodology has shown success in diverse optimization tasks such as chemical reactions 2, 6, 1,functional molecules 7 or nanocrystal shape8. Iterating through this learning cycle has recently enabled the rapid identification of optimal conditions within extensive search spaces 9, 10.
@@
While canonical BO has recently been popularized among experimentalists, the experiment design may benefit if the practitioner can collect data at different degrees of reliability while paying a lower price. Additional experimental evidence that may be readily available can be integrated within the model representing the process to be optimized (e.g., low-precision experiments conducted with bench-top nuclear magnetic resonance can be integrated into more expensive, high-precision experiments 11). In classical experimental design, the inclusion of such information sources representing different reliabilities is referred to as multi-fidelity Bayesian Optimization (MFBO). Within this setting, a specific cost is assigned to each information source — hereafter defined as fidelity. The multi-fidelity probabilistic model learns the process of interest by extracting knowledge from data available at different fidelities and understanding their interplay. The policy for querying new experimental data at a given fidelity also takes into account the overall cost. By combining low-fidelity (LF) and high-fidelity (HF) points, the overall optimization cost can be reduced compared to the single-fidelity BO (SFBO). Figure 1 exemplifies how the iterative cycle of MFBO may reduce the overall cost, compared to the standard SFBO.
@@
Although MFBO has garnered interest in the past decade within the ML community, leading to a myriad of model definitions and acquisition policies 12, 13, 14, 15, researchers have only lately started to integrate them within their design methodologies. In the chemical domain, there has been recent interest in incorporating cost awareness into the BO loop16, 17. Regarding the specific multi-fidelity approach, several studies have successfully applied MFBO to the materials discovery domain 18, 19, 20, 21, 22, 23, drug discovery24 or reactor design25. However, there is a fundamental lack in the assessment of MFBO performance, and due to this, each work represents MFBO performance in different ways 26, 19, 22, 23. Although some metrics have been proposed 18, 22, the lack of clarity and unified criteria on how to assess the benefits of MFBO in the chemical domain has hindered its widespread adoption among the experimental community27, 28. Several factors come into play when assessing its impact, yet the application of MFBO can also be detrimental to the overall optimization process 29. It has been also shown how in the long run the advantage of MFBO over SFBO can be lost 27. Therefore, it is crucial to unify and provide a reliable method to decide when MFBO is better than SFBO.
@@
In this work, we propose a series of guidelines for when to use MFBO within the experimental design pipeline. We conduct an exhaustive experimental investigation on both standard MFBO problems — ubiquitous in BO literature — as well as chemistry-based ones. Initially, we optimize two synthetic problems to assess the behavior of a multi-fidelity model in simulated black-box scenarios, detecting unfavorable situations where the application of MFBO does not offer an advantage over SFBO. We exhaustively scan MFBO experimental parameters (namely, cost ratio and informativeness of the LF source) to identify trends that indicate promising experimental settings for when to apply MFBO. We then progress to chemistry-based problems, tackling three real challenges in molecular and materials optimization, where MFBO successfully outperforms SFBO. Finally, we extract guidelines and recommendations based on the synthetic and real benchmarks to inform prospective users on which scenarios can favor MFBO over SFBO. This work offers a comprehensive view and a reference guide for applying MFBO in the molecules and materials discovery domain.
@@
Results
Shedding light into MFBO failure modes
We test the performance of MFBO methods over their SFBO counterparts on two different synthetic functions, Branin and Park (with 2 and 4 dimensions respectively), on two different scenarios (favorable and unfavorable). Synthetic functions are commonly employed as black-box problems to test the performance of BO algorithms (see Synthetic functions). In the case of MFBO, the output of the functions can be biased using a parameter $\alpha$ to provide lower accuracy information sources. This parameter is related to the informativeness of the LF source (in our case, we quantify the informativeness of the LF by computing the $R^{2}$ with respect to the HF, see Metrics). For the favorable scenario, we tune the $\alpha$ parameter to provide a highly informative LF source (LF $R^{2}>0.9$). We set the cost of querying a HF point to 1 and the cost of a LF point to 0.1 ($\rho$ = 0.1). Figure 2 shows the results for the favorable scenario. In this case, both MFBO methods outperform their SFBO counterparts, with maximum discounts of 0.53 and 0.33 for Branin and Park respectively. The MFBO runs get to lower regrets quicker, spending less resources by exploiting the use of LF and HF sources of information. The resulting optimization runs are therefore desirable as they effectively leverage the access to the LF to guide the optimization of the HF level.
Although under the previous settings MFBO provides a higher performance than SFBO, a change in the LF conditions dramatically affects the result of the optimization. In the unfavorable scenario, we set the cost of the LF source to 0.5 (only half of the HF cost, $\rho$ = 0.5) and decrease the informativeness of the functions ($R^{2}$ $<$ 0.75). In this case, the performance of both methods decreases with respect to the favorable scenario. In Branin, the maximum $\Delta$ drops to -0.07, whereas in Park it drops to a maximum $\Delta$ of -0.11. In both cases the favorable trend is reverted, and MFBO loses its advantage over SFBO. These results illustrate how MFBO performance is affected by the cost and the informativeness of the LF source, losing its advantage over SFBO if the LF source is not informative and cheap enough. It also reflects how the MFBO setting is problem-dependent and it may be more robust towards problem changing situation
Apart from problem conditions, model choice may also affect MFBO performance. Previous studies have proposed and used many MF models, mainly GPs refs but also DNNs13, 30. We compared the standard BoTorch multi-fidelity model with a MultiTask GP12 to see how model choice can affect the results. The main difference between the two models lies in how fidelities are encoded in the GP. SI section A shows a preliminary study on the fidelity kernel and how both models offer similar performances in regression and MFBO tasks with the Branin function (figures 7 and 8 respectively). In addition, using two independent Gaussian Processes (GPs) for high-fidelity (HF) and low-fidelity (LF) data would fail to model the correlation between them, as each GP would learn its own independent function without leveraging shared structure. The key advantage of a multi-fidelity GP is that it introduces a covariance structure between HF and LF data with the kernel $k_{I}S$. This correlation enables the model to transfer information across fidelities, leading to more data-efficient learning and better uncertainty quantification, especially when HF data is scarce. As an example, the regression task in the Supplementary Information A.6 shows how the incorporation of lower fidelities improves the uncertainty estimation at the HF level. Therefore, we stick to the use of the default model provided in BOTorch, focusing on the external factors that can affect the experiments.
Figure 2: MFBO results are dependent on problem conditions. In the favorable scenario (highly informative and cheap LF source, $\rho$ = 0.1, $R^{2}$ > 0.9), MFBO offers better performance than SFBO on the synthetic functions (maximum $\Delta$ of 0.53 and 0.33 for Branin and Park, respectively). The optimization traces show how lower regrets are found at lower budgets under this scenario. When the conditions are changed and a low informative and more expensive LF is used ($\rho$ = 0.5, $R^{2}$ < 0.75), MFBO performance is degraded. Under this scenario, the optimization traces are similar, and MFBO performance is notably reduced (minimum $\Delta$ of -0.07 and -0.11 for Branin and Park, respectively). The reference budget unit cost corresponds to a single HF evaluation.
Finding suitable scenarios for MFBO
We investigate in detail the effect of the LF cost and informativeness on MFBO performance (measured by $\Delta$, see Metrics) for the two previous synthetic functions. We run the MFBO method with different combinations of LF cost and informativeness and compare it to SFBO. Figure 3 shows the computed $\Delta$ for each run as a result of the previous parameters. The $x$ axis of the heatmap represents the $\alpha$ value that was used to bias the synthetic function, and the upper plot shows the computed $R^{2}$ for each value. In both cases, a gradient can be observed where the progression towards cheaper and more informative LF sources provides better MFBO performances (higher $\Delta$). In the case of Branin, the lower values of $R^{2}$ in the synthetic functions are associated with a less informative LF approximation, which translates to lower values of $\Delta$. In the same way, in Park $\Delta$ is also decreased when the $R^{2}$ is lower, giving negative results when its value is close to 0. This result is consistent as a more inaccurate source of information is likely to reduce the performance of the MFBO method. In terms of cost, there is also an inverse correlation in both cases, where lower $\rho$ (cheaper LF sources) generate higher discounts. This is an expected trend, as the model can access the approximate sources at a lower cost, allowing a wider exploration of the problem space for the same price.
Although $\Delta$ is reported at a value of $\tau$ = 0.9 for all the cases, we also investigated the effect of $\tau$ on the final discount. In previous works, the MFBO performance was reported to be lower than SFBO at smaller budgets23, and the authors noted that in the long-run MFBO may also lose its advantage over SFBO27. This trend was observed in both functions and AFs, where low values of $\tau$ and a $\tau$ of 1 provided negative discounts (see figure 10 in the SI). Although the temporal dependence of the MFBO is hard to control for the experimenter due to the uncertainty of when to stop the sampling, the previous figure shows how there exist "sweet spots" where MFBO can exploit the cheap LF source with the available budget to provide an advantage over SFBO.
The results follow the same trend in both cases independently of the acquisition function. Although there are differences in the absolute $\Delta$ values between the two problems (due to the specific landscape of each synthetic function), the regions of high performance are localized in the same areas of the heatmaps. It is worth noticing that in Branin-2D, only $\rho$ values below 0.1 provide the best MFBO performance, similar to what was found in the synthetic functions experiment of a similar study28, and in previous studies on materials discovery23. These results provide a qualitative understanding of the effect of cost and informativeness on the performance of MFBO. The general high-performance region is located around cheap and highly informative LF sources, and a good experimental setup for MFBO application should aim to find suitable LF sources.
Application of MFBO to chemistry and materials problems
We evaluate three benchmarks in chemistry and materials science to study the performance of MFBO in more realistic experimental scenarios . The tasks correspond to scenarios where a cheap and informative LF source is available (see Methods for a detailed explanation of each benchmark). Figure 4a illustrates the results of MFBO and SFBO in these scenarios. In all cases, the MFBO method can get to lower regrets at a lower cost than the SFBO by leveraging the cheaper source of information at hand. In the COFs benchmark (problem settings: $\rho$ = 0.065, $R^{2}$ = 0.98), the MFBO gets to regrets close to 0 in an early phase of the optimization, whereas the SFBO only does it at the end of the optimization (the maximum $\Delta$ is 0.68). In the solvation energy benchmark (problem settings: $\rho$ = 0.1, $R^{2}$ = 0.88), MFBO also gets to regrets close to 0 at a lower budget than SFBO. This example also integrates computed values with experimental measurements to reduce the spent budget. In the polarizability benchmark (problem settings: $\rho$ = 0.167, $R^{2}$ = 0.99), the MFBO provides a maximum discount of 0.56 over the SFBO. In this case, the use of a cheap Hartree-Fock computation is successfully combined with the experimental measurements to find the molecule with the highest polarizability. Figure 4b shows the proportion of queries to each fidelity level in the different tasks. By distributing the available budget between the HF and LF levels, the MFBO methods can integrate the cheaper information of the LF approximations to improve the optimization results. The number of calls to each query depends on the specific acquisition function and task. However, in all cases the proportion of calls to the HF level is lower than 0.4, suggesting that this proportion can provide a threshold for a successful MFBO optimization. A previous study suggested a proportion of LF-HF points of 5:1 as the optimal case for good MFBO performance23, which is similar to the 4:1 ratio observed in the COFs and polarizability benchmarks. In these real scenarios, the MFBO methods can exploit the low accuracy but cheap LF levels by querying them more often than the HF level.
Figure 4: Molecules and materials benchmark results. a) Optimization performance for the MFBO and the SFBO methods in each benchmark (see Methods for a description of each task). The maximum $\Delta$ in each case is 0.68, 0.59 and 0.56 for the COFs, solvation energy and polarizability tasks, respectively. In all the cases, MFBO is able to achieve lower regrets than SFBO by spending a lower budget. b) Proportion of queries to each fidelity source in the MFBO setting. The MFBO method balances information gain and budget spent by querying more often the LF values (between 0.6 and 0.8 in the ratio of queries) while also using information from HF values. c) Experimental negative example. Molecule polarizability benchmark conditions were changed so the LF source was expensive and less informative. The resulting $\Delta$ is 0.07 and -1 for the MES and EI families respectively, showing how unfavorable conditions can override the MFBO advantage over SFBO in a real experimental setting. The reference budget unit cost corresponds to a single HF evaluation in all cases.
We also include a negative example where the conditions of one of these benchmarks are changed to an unfavorable scenario and the associated original MFBO performance is overridden. We artificially change the LF conditions of the molecule polarizability benchmark ($\rho$ = 0.5 and $R^{2}$ = 0.49) to illustrate how $\Delta$ is reduced with respect to the previous case. Figure 5c shows the result for the optimization, where $\Delta$ is reduced from a maximum value of 0.56 to -1 in the EI case (that is, the MFBO gives worse result than the SFBO), providing no advantage over SFBO if compared with the previous favorable scenario (similar to the results described for the synthetic functions, see Results). This highlights the importance of using adequate LF sources, and how their incorporation can be detrimental otherwise29. The previous benchmarks show how the application of MFBO in suitable scenarios provides an advantage over standard SFBO cases and effectively decreases the optimization cost. By locating experiments where the cost and informativeness of the LF source are adequate, we can successfully translate the advantages of MFBO into a realistic experimental task.
Guidelines for MFBO in chemistry and materials
We discuss and provide some recommendations to run MFBO in an experimental binary setting from the results obtained in this study. Our results highlight how the focus should be directed to the cost and informativeness of the LF source. Figure 5 displays a guiding flowchart following the previous considerations. Using these conditions does not guarantee superior performance over SFBO, but provides a guiding principle for the successful application of MFBO in real scenarios.
In general, a user should first verify that the cost of the auxiliary experiment is cheap enough to run MFBO. All experimental benchmarks considered in this study employed a cost ratio below 0.167, with the synthetic benchmarks achieving optimal performance at $\rho=0.05$. Notably, this observation aligns with prior findings in bandgap optimization, where the most effective MFBO performance was reported at a cost ratio of 5% 23. This suggests that even though more expensive LF sources may also provide advantages, $\rho$ values around 0.05 can be a good starting point to try MFBO. In the same way, the user should check and estimate the degree of informativeness of the auxiliary experiment to decide if it provides a reasonable approximation of the real problem. The synthetic benchmark showed how uncorrelated ($R^{2}\leq 0$) LF functions always gave negative MFBO performance. In the experimental benchmark, all the LF $R^{2}$ values were above 0.85, and when artificially decreasing it to 0.49 and increasing $\rho$ from 0.167 to 0.5 in the polarizability case, the MFBO performance degraded. Previous work used LF values where $R^{2}$ = 0.8423. Therefore we recommend $R^{2}$ values over 0.80 as a guiding principle to use informative LF sources.
It is important to notice that informativeness measurement is difficult to estimate beforehand. In this case, we propose $R^{2}$ as a guiding metric, but other informativeness measurements could be employed (for instance, metrics that can capture nonlinear trends). Other studies have used the Pearson correlation coefficient28 or the Matthew correlation coefficient 24 to estimate the informativeness of the LF sources. Also, the cost is arbitrary depending on the user and experiment, as different setups may prioritize some costs over others. For example, experiment time is the most intuitive cost measurement, but also experiment price or any other relevant quantity could be used. In addition, it may be difficult to accurately estimate these parameters when starting the optimization from scratch. A potential option in a realistic scenario is to use the initial sampled points to evaluate the cost and informativeness of the LF source with respect to the sampled MF by running a simple regression task to check if the MF surrogate improves the SF predictions. This alternative estimation is left for future work.
Despite all these challenges, the idea holds: the user should reflect if the conditions are met for MFBO to avoid unreliable LF sources29 or long-term performance degradation of the MFBO method27 in order to successfully apply MFBO in real scenarios.
Figure 5: Guidelines for MFBO application in chemistry and materials. Given a potential optimization problem where a cheaper source of information is available, the user should estimate if this source is cheap enough (e.g. $\rho$<0.1, based on this work) and informative enough (e.g. $R^{2}$>0.8, based on this work) before running MFBO . These parameters could be estimated on the initial data sampling or using previous experiments. If these conditions are not met, they should use SFBO otherwise.
Discussion
Multi-fidelity Bayesian optimization (MFBO) offers a promising approach to reducing costs in optimization by leveraging inexpensive, approximate sources of information. In the chemical sciences, MFBO can accelerate optimization while maintaining cost efficiency. However, its utility is not universal; in some cases, it may prove less effective or even counterproductive29. To address these concerns, we conducted a comprehensive study to determine under which conditions MFBO outperforms standard single-fidelity Bayesian optimization (SFBO).
Building on previous work in the field, we introduce two key metrics for comparison: a standardized regret, $r$, and a discount metric, $\Delta$ (see Metrics). Our analysis of synthetic benchmarks demonstrates that MFBO performance is highly dependent on the characteristics of the low-fidelity (LF) source. Specifically, not all LF conditions are conducive to the effective use of MFBO. To systematically identify favorable conditions, we performed a grid-based exploration of LF cost and informativeness against the discount metric, $\Delta$. The results indicate that both $\rho$ and $R^{2}$ significantly influence the discount, with the most favorable outcomes achieved when the LF source is both highly informative and low-cost. The method was validated using three real-world experimental benchmarks in molecular and materials discovery, demonstrating that MFBO consistently outperforms SFBO in reducing optimization costs, provided the LF source is sufficiently inexpensive and informative. Based on these findings, we developed a flowchart to guide the application of MFBO (see Figure 4).
Our results resonate with previous studies suggesting a favorable cost ratio of 0.0523 and showing how informative sources are required for MFBO to outperform SFBO28. In addition, the optimal LF/HF query proportion of 4:1 observed in the experimental benchmarks also aligns with previous observations23 that found 5:1 as the best ratio, although we use an adaptive acquisition function instead of a greedy heuristic. Although the modelling was not the main focus of the study, other works used different surrogates like one-hot encoded GP22 or a random forest23, and comparing the MFBO and SFBO performance of all the surrogates could be a future research direction to complement the guidelines. In addition, model uncertainty calibration may also play a key role during the acquisition stage, and studying its effect on MFBO performance would help to clarify how its effect interacts with poor or high LF informativeness.
Our work provides a structured decision-making framework for determining the applicability of MFBO. While our analysis focused on a specific surrogate model and two acquisition functions, the approach is extensible to other surrogate models and acquisition function families. Future research will expand this framework to include models such as Bayesian neural networks (BNNs)31, and explore additional experimental applications, evaluating the impact of varied feature spaces or additional fidelity levels. This study offers valuable insights for practitioners seeking to integrate MFBO into their experimental workflows, paving the way for more routine application of MFBO in chemical and materials optimization.
@@ Methods
where f is the target black-box function and $\mathcal{X}$ is the space of all possible input candidates. Considering $\mathcal{D}=[(x_{1},y_{1}),(x_{2},y_{2}),...(x_{n},y_{n})]$ a dataset containing pairs of inputs-outputs $({x}_{n},y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ of our problem, we can use a surrogate model to learn the relationships between the data in the form $y_{i}=f({x}_{i})+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. The surrogate is commonly a Gaussian Process (GP), a non-parametric model that provides uncertainty quantification 32. GPs are usually defined as $f(x)\sim\mathcal{GP}(0,k({x},{x^{\prime}}))$, placing a zero-mean prior and a covariance given by a kernel function $cov[f({x}),f({x^{\prime}})]=k({x},{x^{\prime}})$ that measures the similarity between inputs. Given a dataset $\mathcal{D}$, the predictive mean and variance are given by the following expressions
@@
where ${k_{x}}=[k({x},{x_{1}}),...,k({x},{x_{n}})]^{\top}\in\mathbb{R}^{n}$, $K\in\mathbb{R}^{n\times n}=(k({x_{i}},{x_{j}}))_{1\leq i,j\leq n}$, ${y}=[y_{1},...,y_{n}]$, I is the $n\times n$ identity matrix, and $\sigma^{2}_{\epsilon}$ is the variance of the observation noise. In order to optimize the target problem, after training the surrogate with the available data, a decision is made to select the next points by maximizing the acquisition function. Acquisition functions are heuristics that decide the next point to query by balancing exploration and exploitation. Exploration promotes sampling in regions of high uncertainty to improve the model’s understanding of the search space, while exploitation prioritizes regions near the current best observations to refine promising solutions. Each acquisition function family offers different tradeoffs between these two terms. The next inputs are given by
$x_{n+1}=\operatorname*{argmax}_{{x}\in\mathcal{X}}\texttt{acqf}({x;\mathcal{D}})$
where acqf is the acquisition function, which uses the surrogate trained on the sampled dataset $\mathcal{D}$ to select the next query among the candidates. Common acquisition functions include Expected Improvement (EI)33, Maximum Entropy Search (MES)34 and Knowledge Gradient (KG)35.
Our multi-fidelity BO extends the previous setup by considering an extra input parameter for the surrogate model, the fidelity l. This parameter indicates which source of information the other inputs correspond to (e.g., high fidelity, HF, or low fidelity, LF). In our case, we use a discrete fidelity setting, where $l\in[l_{0},l_{1},...l_{m}]$ and $l_{m}$ corresponds to the highest fidelity among the $m$ levels. In all cases, we limit the analysis to two levels of fidelity ($l_{0}$ and $l_{1}$ and $m=1$) , but the methodology can be extended to more levels (see Supplementary Information A.6 for an example of a 3-fidelities regression task). Additionally, these models assume that the cost only depends on the fidelity level $l$ and not on the search space $x$. We employed a modified GP proposed in 14 that extends the previously described surrogate modelling to the multi-fidelity space of $\mathbb{R}^{n+1}$ dimensions, chosen due to its popularity amongst MFBO applications14, 19, 29. The modified kernel is defined as $k(({x},l),({x^{\prime}},l^{\prime}))$, where fidelity is introduced by defining a separate kernel to model the input space, and a kernel to model the correlation and interaction between the different fidelity levels. The mathematical expression is:
@@
Given a dataset of inputs with their associated fidelities and outputs, the model can be trained using standard maximum marginal log-likelihood optimization to obtain the optimal hyperparameters $\lambda_{i}$, $c$ and $\delta$. In the experiments, we set the $l$ values of the HF and LF levels to $\frac{2}{3}$ and $\frac{1}{3}$ respectively following previous studies that used this kernel in a binary setting19. The choice of $l_{0}$, and $l_{1}$ values was done to ensure the c and the exponents of the $k_{IS}$ have an impact on the kernel value, which may not be the case if other pair values of l and l’ were chosen such as 0/1.
@@
$x_{n+1},l_{n+1}=\operatorname*{argmax}_{(x,l)\in\mathcal{X}\times[m]}\texttt{acqf}(x,l;\mathcal{D})$
In this setting, the acquisition function selects the input-fidelity pair by also considering the cost of each fidelity. In the experiments, we use two families of acquisition functions, with their single and multi-fidelity versions: Expected Improvement (EI)36 and Maximum Entropy Search (MES)15.
The expression of the multi-fidelity EI function is
[…]
Where $\mathcal{D}_{n}$ is the dataset acquired at a step n of the optimization. The first term is the EI in the objective property of the HF simulation, and balances exploration and exploitation by favoring candidates with high and/or uncertain predicted objective. This term is computed as the expectation of the predicted HF values $Y^{m}(\mathbf{x})\Big{|}\mathcal{D}_{[n]}$ greater than the best HF value observed so far, $\hat{y}^{m*}_{[n]}$. The second term is the correlation between the target property of x under the l fidelity and the high-fidelity $l_{m}$, which penalizes uncorrelated low-fidelity values. The last term is the inverse of the cost ratio between the high fidelity source $l_{m}$ and the fidelity l, which rewards cheaper samples. The composed function rewards candidates that have a high utility per cost for the target objective.
The expression of the multi-fidelity MES function is
$x_{n+1},l_{n+1}=\operatorname*{argmax}_{(x,l)\in\mathcal{X}\times[m]}\left(\mathbb{H}(y_{m}\mid x,\mathcal{D}_{n})-\mathbb{E}_{f^{*}\mid x,\mathbb{D}}[\mathbb{H}(y_{m}\mid f^{*},x,\mathcal{D}_{n})]\right)\times\frac{1}{\rho}$
Where $\mathbb{H}$ is the entropy, which measures uncertainty or the information content of a random variable or distribution. The first term $\mathbb{H}(y_{m}|x,\mathcal{D}_{n})$ represents the entropy of the function value at the highest fidelity level $l_{m}$. It reflects the uncertainty about the objective function at the candidate point x before making a new observation. The second term $\mathbb{E}_{f^{*}\mid x,\mathbb{D}}[\mathbb{H}(y_{m}\mid f^{*},x,\mathcal{D}_{n})]$ denotes the expected entropy of the posterior distribution of the highest fidelity function value at point x, conditioned on a hypothesized realization of the global optimum $f^{*}$. This conditional entropy indicates the uncertainty after observing the hypothetical optimal function value. The difference between these two entropy terms is known as the information gain, which measures how much observing a particular point at fidelity l is expected to reduce uncertainty about the optimal solution. The third term is the inverse of the cost ratio between the high fidelity source and the fidelity l. MES rewards candidates that bring a reduction of high-fidelity entropy scaled by their respective cost.
We also initially tested Knowledge Gradient (KG)14, but we discarded it due to its computational expense.
In all cases, the experiments are run for 20 independent seeds. We assign 10% of the total (pre-selected) optimization cost to the initial sampling, which we refer to as sampling cost (the unit of cost in this work is a HF evaluation). In all cases, the selected budget value is selected to achieve low regrets in the SFBO case. Then, we assign 50% of this budget to HF samples and 50% to low fidelity samples (the number of HF and LF samples depend on the relative cost of each one). Given that we define cost in terms of HF evaluation units, if a sampling cost of 4 is used, we take 2 HF samples and 2/$\rho$ LF samples. For continuous spaces, we employ a Latin Hypercube Sampling (LHS) strategy. In the case of categorical benchmarks, we adopt a methodology similar to the one described in previous work19. This approach involves the random selection of an initial example, followed by the sequential selection of the most spatially distant samples relative to those already chosen (Furthest Point sampling).
In the synthetic functions, the total budget corresponds to 50 HF queries. In the chemistry and materials benchmarks, the total budget corresponds to 30 HF queries for the COFs and polarizability, and 50 HF for the solvation energy task. Data collection is sequential. All the experiments are run using BOTorch 37 as a standard Bayesian Optimization framework.